OpenStudy (anonymous):

Which of the ordered pairs in the form (x, y) is a solution of this equation? 3x-y/4 = 11 (3, –8), (4, 4) A. The first is not a solution, but the second is. B. The first is a solution, but the second is not. C. Neither is a solution. D. Both are solutions.

OpenStudy (anonymous):

@Nurali ?

OpenStudy (nurali):

3x-y/4 = 11 (3, –8) 3.3+8/4=11 9+8/4=11 9+2=11 11=11 3x-y/4 = 11 (4, 4) 3.4-4/4=11 12-1=11 11=11 Both are solutions.

OpenStudy (anonymous):

Thank you!! <3

OpenStudy (nurali):

My Pleasure.